# ziji: a model to calculate X-ray spectra of black hole X-ray binaries

### v1

@@ Abstract
In the disk-corona model, the X-ray spectrum of a stellar-mass black hole in an X-ray binary is characterized by three components: a thermal component from a thin and cold accretion disk, a Comptonized component from a hot corona, and a reflection component produced by illumination of the cold disk by the hot corona. In this paper, we present ziji, which is a model to calculate the X-ray spectrum of black hole X-ray binaries in the disk-corona model. The accretion disk is described by the Novikov-Thorne model. The reflection spectrum is produced by the direct radiation from the corona as well as by the returning radiation of the thermal and reflection components and is calculated considering the actual spectrum illuminating the disk. If we turn the corona off, the reflection spectrum is completely generated by the returning radiation of the thermal component, as it may happen for some sources in soft spectral states. The user decides the radial density profile of the accretion disk and the ionization parameter is calculated self-consistently at any radial coordinate of the disk from the illuminating X-ray flux and the local electron density. We show the predictions of ziji in different regimes. We discuss current limitations of our model as well as the next steps to improve it.
@@ 1. Introduction
Figure 1.— Disk-corona model. Figure from Bambi (2022a) under the terms of the Creative Commons Attribution 4.0 International License.
@@
There is a relatively rich literature on the study of the simplifications in current reflection models and on their impact on current measurements of accreting black holes, as well as attempts to remove such simplifications. Simplifications in current reflection models mainly appear in the structure of the accretion disk, in the description of the corona, and in the calculations of the reflection spectrum in the rest-frame of the material in the disk. Even some relativistic effects are not properly taken into account. For a non-complete list of these studies, see, for instance, Niedźwiecki & Życki (2008); Wilkins & Fabian (2011, 2012); Niedźwiecki et al. (2016); Gonzalez et al. (2017); Niedźwiecki & Zdziarski (2018); Cárdenas-Avendaño et al. (2020); Riaz et al. (2020, 2021); Abdikamalov et al. (2020, 2021); Shashank et al. (2022); Mall et al. (2022); Mirzaev et al. (2024).
@@
In Mirzaev et al. (2024), we showed that the returning radiation can have a strong impact on the reflection spectrum for some systems and its effect cannot be reabsorbed by a modified emissivity profile with respect to the one produced by the direct radiation from the corona, because the spectrum from the corona is approximately a power law with a high-energy cutoff while the spectrum of the returning radiation is a reflection spectrum. The effect is particularly strong when the inner edge of the disk is very close to the black hole, the corona illuminates mainly the inner part of the accretion disk, and in the case of a low or moderate value of the ionization parameter. For such systems, it is crucial to calculate the reflection spectrum at every radial coordinate of the disk by using the actual spectrum of the incident radiation. If we do not do so, the model cannot fit the data well and some parameters can be significantly overestimated or underestimated (Mirzaev et al., 2024).
@@
In this manuscript, we extend the work in Mirzaev et al. (2024) and we present the model ziji. ziji calculates the thermal spectrum of the cold disk, the Comptonized spectrum from the corona, and the reflection spectrum from the disk. The reflection spectrum is produced by the direct radiation from the corona as well as by the returning radiation of the thermal and reflection components and is calculated considering the actual spectrum illuminating the disk. In Mirzaev et al. (2024), we ignored the thermal spectrum of the disk and therefore there was no returning radiation of the thermal component in the incident spectrum generating the reflection spectrum. We note that the new calculations presented here can be important only for stellar-mass black holes in X-ray binaries, as their thermal spectrum can be peaked in the soft X-ray band and it can thus have an impact on the reflection spectrum in the X-ray band. In the case of supermassive black holes in active galactic nuclei, the thermal spectrum is peaked in the UV band and its returning radiation cannot appreciably affect the reflection spectrum in the X-ray band. We also note that there is some observational evidence that the reflection spectrum can be produced mainly by the returning radiation of the thermal component when a source is in the soft spectral state, where the total spectrum is dominated by the thermal component and the coronal spectrum is weak or absent (Connors et al., 2021).
@@ 2. Model
The accretion disk is described by the Novikov-Thorne model (Novikov & Thorne, 1973; Page & Thorne, 1974). The disk is infinitesimally thin, perpendicular to the black hole spin axis, and the inner edge of the disk is at the innermost stable circular orbit (ISCO) or at a larger radius. The material in the disk moves on nearly-geodesic equatorial circular orbits. From the conservation of mass, energy, and angular momentum, the model predicts the time-averaged radial structure of the disk. The time-averaged energy flux emitted from the surface of the disk at the radial coordinate $r$ is
@@
where in the expression of $\Omega$ the sign $+$ is for corotating orbits and the sign -- is for counterrotating orbits.
@@
As default coronal geometry, we use the lamppost model, where the corona is a point-like source with isotropic emission along the black hole spin axis and is completely specified by the coronal height $h$. This is certainly the simplest choice. However, it is straightforward to consider more complicated coronal geometries with non-isotropic emission (Riaz et al., 2022). A ray-tracing code calculates how the corona (direct radiation from the corona) and the disk (thermal radiation of the disk returning to the disk) illuminate the disk (Mirzaev et al., 2024). The latter is characterized by an electron density profile $n_{\rm e}=n_{\rm e}(r)$ (which can be specified by the user, but in the next section we will consider the simplest case of constant electron density) and its iron abundance $A_{\rm Fe}$. The ionization parameter of the disk $\xi$ at the radial coordinate $r$ is determined by the total incident photon flux $\Phi_{X}$ and the electron density at that radial coordinate
$\xi=\frac{4\pi\Phi_{X}}{n_{\rm e}}\,.$ (9)
For the calculation of the reflection spectrum in the rest-frame of the material in the disk, we use the reflionx code (Ross & Fabian, 2005). We do not use the reflionx table because we want to calculate the reflection radiation with the correct spectrum of the incident radiation, while in the reflionx table it is assumed that the radiation illuminating the disk has a power law spectrum (which is a good approximation only for the direct radiation from the corona). After calculating the reflection spectrum from the direct radiation from the corona and the returning radiation of the thermal spectrum, the model includes the effect of the returning radiation of the reflection spectrum itself with a few iterations. The details of these calculations have been already presented in Mirzaev et al. (2024). At the end, we have a thermal component from the Novikov-Thorne accretion disk, a Comptonized component from the corona, and a reflection component calculated with the actual spectrum of the radiation illuminating the disk (direct radiation from the corona and returning radiation from the disk). Table 1 lists the parameters of the model in its simplest version.
@@ 3. Results
In this section, we present the predictions of ziji. We set the black hole mass to $10~{}M_{\odot}$. To maximize the impact of the returning radiation, we assume that the black hole spin parameter is $a_{*}=0.998$ and that the inner edge of the accretion disk is at the ISCO radius, $r_{\rm in}=r_{\rm ISCO}$. There is no emission from the plunging region $r<r_{\rm in}$. As shown in Fig. 1 in Mirzaev et al. (2024), a significant fraction of the radiation emitted by the disk returns to the disk only in the very inner part of the accretion disk, when $r<2~{}r_{\rm g}$, where $r_{\rm g}=M$ is the black hole gravitational radius. Already at $r=2~{}r_{\rm g}$, most of the emitted radiation can escape to infinity and this fraction approaches 1 as the emission radius increases. We assume a lamppost corona and we set $h=5~{}r_{\rm g}$, where $h$ is the coronal height (i.e., the radial coordinate in Boyer-Lindquist coordinates of the point-like corona along the black hole spin axis).
In Fig. 2, we show the spectra in the rest-frame of the material in the disk at the radial coordinates $r=1.3~{}r_{\rm g}$ (left panels), $5~{}r_{\rm g}$ (central panels), and $10~{}r_{\rm g}$ (right panels) when the Eddington-scaled luminosity of the source is $\lambda=0.01$. The top panels are for $\rho=100$ (the spectrum of the source is essentially the Comptonized spectrum from the corona), the central panels are for $\rho=1$ (the corona and the disk have the same luminosity), and the bottom panels are for $\rho=0$ (there is no corona and the reflection spectrum is generated exclusively by the returning radiation of the thermal spectrum of the disk). The spectrum of the corona is always described by a power law with photon index $\Gamma=2$ and a high-energy cutoff $E_{\rm cut}=300$ keV (in the rest-frame of the corona). The electron density of the disk is $n_{\rm e}=10^{22}$ cm-3 at every radial coordinate and the ionization parameter $\xi$ is calculated from Eq. (9). The iron abundance is $A_{\rm Fe}=1$. In every panel, we show the spectrum of the direct radiation from the corona (yellow solid line), the spectrum of the returning radiation of the thermal component (black solid line), the reflection spectrum generated by these two components (blue dashed line), and the final reflection spectrum that includes even the returning radiation of the reflection spectrum (red dashed line).
Fig. 3 is like Fig. 2 but for $\lambda=0.1$. The higher luminosity of the source implies a higher ionization parameter and the spectra in the rest-frame of the material in the disk show fewer emission lines if compared to the panels in Fig. 2 with the same radial coordinate $r$. This is in particular the case for the panels for $r=1.3$ $r_{\rm g}$, because the ionization parameter is too high.
Fig. 4 shows the reflection spectra of the whole disk as observed far from the source. The luminosity of the source is $\lambda=0.01$ (top panels) and 0.1 (bottom panels). The left panels are for $\rho=100$, the central panels are for $\rho=1$, and the right panels are for $\rho=0$.
@@
Figure 2.— Synthetic reflection spectra in the rest-frame of the material in the disk at different radial coordinates ($r=1.3$, 3, and 10 $r_{\rm g}$) when the Eddington-scaled luminosity of the source is $\lambda=0.01$. In the top panels, the spectrum of the source is dominated by the spectrum from the corona ($\rho=100$). In the central panels, the luminosity of the corona and of the disk are equivalent ($\rho=1$). In the bottom panels, there is no corona and the reflection spectrum is entirely generated by the returning radiation of the thermal component ($\rho=0$). In every panel, yellow solid lines are for the direct incident spectra from the corona, black solid lines are for the spectra of the returning radiation of the thermal component, blue dashed lines are for the reflection spectra produced by the direct radiation from the corona and the returning radiation of the thermal component, and red dashed lines are for the final reflection spectra including the returning radiation of the reflection radiation. See the text for more details.
@@
Figure 4.— Synthetic reflection spectra of the whole disk as observed far from the source. In the top panels, the Eddington-scaled luminosity of the source is $\lambda=0.01$. In the bottom panels, we have $\lambda=0.1$. In the left panels, the luminosity of the source is dominated by the corona ($\rho=100$). In the right panels, there is no corona and therefore the reflection spectrum is produced by the returning radiation of the thermal component ($\rho=0$). In the central panels, the luminosity of the corona and of the disk are equivalent ($\rho=1$). See the text for more details.

### v2

# X-ray spectra of black hole X-ray binaries with returning radiation

@@ Abstract
In the disk-corona model, the X-ray spectrum of a stellar-mass black hole in an X-ray binary is characterized by three components: a thermal component from a thin and cold accretion disk, a Comptonized component from a hot corona, and a reflection component produced by illumination of the cold disk by the hot corona. In this paper, we assume a lamppost corona and we improve previous calculations of the X-ray spectrum of black hole X-ray binaries. The reflection spectrum is produced by the direct radiation from the corona as well as by the returning radiation of the thermal and reflection components and is calculated considering the actual spectrum illuminating the disk. If we turn the corona off, the reflection spectrum is completely generated by the returning radiation of the thermal component, as it may happen for some sources in soft spectral states. After choosing the radial density profile of the accretion disk, the ionization parameter is calculated self-consistently at any radial coordinate of the disk from the illuminating X-ray flux and the local electron density. We show the predictions of our model in different regimes and we discuss its current limitations as well as the next steps to improve it.
@@ 1. Introduction
Figure 1.— Disk-corona model. Figure from Bambi (2022) under the terms of the Creative Commons Attribution 4.0 International License.
@@
There is a relatively rich literature on the study of the simplifications in current reflection models and on their impact on current measurements of accreting black holes, as well as attempts to remove such simplifications. Simplifications in current reflection models mainly appear in the structure of the accretion disk, in the description of the corona, and in the calculations of the reflection spectrum in the rest-frame of the material in the disk. Even some relativistic effects are not properly taken into account. For a non-complete list of these studies, see, for instance, Niedźwiecki & Życki (2008); Wilkins & Fabian (2011, 2012); Niedźwiecki et al. (2016); Gonzalez et al. (2017); Niedźwiecki & Zdziarski (2018); Cárdenas-Avendaño et al. (2020); Riaz et al. (2020, 2021); Abdikamalov et al. (2020, 2021); Shashank et al. (2022); Mall et al. (2022); Mirzaev et al. (2024a).
@@
In Mirzaev et al. (2024a), we showed that the returning radiation can have a strong impact on the reflection spectrum for some systems and its effect cannot be reabsorbed by a modified emissivity profile with respect to the one produced by the direct radiation from the corona, because the spectrum from the corona is approximately a power law with a high-energy cutoff while the spectrum of the returning radiation is a reflection spectrum. The effect is particularly strong when the inner edge of the disk is very close to the black hole, the corona illuminates mainly the inner part of the accretion disk, and in the case of a low or moderate value of the ionization parameter. For such systems, it is crucial to calculate the reflection spectrum at every radial coordinate of the disk by using the actual spectrum of the incident radiation. If we do not do so, the model cannot fit the data well and some parameters can be significantly overestimated or underestimated (Mirzaev et al., 2024a).
@@
In this manuscript, we extend the work in Mirzaev et al. (2024a) and we present the model ziji, which is available on GitHub [^1] https://github.com/ABHModels/ziji/. and Zenodo (Mirzaev et al., 2024b). ziji assumes a lamppost corona and calculates the thermal spectrum of the cold disk, the Comptonized spectrum from the corona, and the reflection spectrum from the disk. The reflection spectrum is produced by the direct radiation from the corona as well as by the returning radiation of the thermal and reflection components and is calculated considering the actual spectrum illuminating the disk. We note that current public models assume that the radiation illuminating the disk is either a power law/Comptonized spectrum (like xillver, García et al., 2013) or a single-temperature blackbody spectrum (like xillverNS, García et al., 2022), but there is no model that considers the possibility that the radiation illuminating the disk is some combination of these two components. Moreover, the returning radiation of the thermal component is not a single-temperature blackbody spectrum but a multi-temperature blackbody spectrum whose exact shape is different at different radii because the returning radiation illuminating a certain point is the combination of the returning radiation from different parts of the accretion disk, with different temperature and different redshift. Our model also includes higher-order reflections, because even the reflection spectrum can return to the disk and produce reflection, and this further changes the actual spectrum of the incident radiation and, in turn, of the reflection component. In Mirzaev et al. (2024a), we ignored the thermal spectrum of the disk and therefore there was no returning radiation of the thermal component in the incident spectrum generating the reflection spectrum. We note that the new calculations presented here can be important only for stellar-mass black holes in X-ray binaries, as their thermal spectrum can be peaked in the soft X-ray band and it can thus have an impact on the reflection spectrum in the X-ray band. In the case of supermassive black holes in active galactic nuclei, the thermal spectrum is peaked in the UV band and its returning radiation cannot appreciably affect the reflection spectrum in the X-ray band. We also note that there is some observational evidence that the reflection spectrum can be produced mainly by the returning radiation of the thermal component when a source is in the soft spectral state, where the total spectrum is dominated by the thermal component and the coronal spectrum is weak or absent (Connors et al., 2021).
@@ 2. Model
The accretion disk is described by the Novikov-Thorne model (Novikov & Thorne, 1973; Page & Thorne, 1974). The disk is infinitesimally thin, perpendicular to the black hole spin axis, and the inner edge of the disk is at the innermost stable circular orbit (ISCO) or at a larger radius. The material in the disk moves on nearly-geodesic equatorial circular orbits. From the conservation of mass, energy, and angular momentum, the model predicts the time-averaged radial structure of the disk. The time-averaged energy flux emitted from the surface of the disk at the radial coordinate $r$ is (Page & Thorne, 1974; Bambi, 2017)
@@
where in the expression of $\Omega$ the sign $+$ is for corotating orbits and the sign -- is for counterrotating orbits. We note that Eq. (2) gives exactly the same temperature profile as Page & Thorne (1974) and can be applied to any stationary, axisymmetric, asymptotically flat, and circular spacetime.
@@
As default coronal geometry, we use the lamppost model, where the corona is a point-like source with isotropic emission along the black hole spin axis and is completely specified by the coronal height $h$. This is certainly the simplest choice. However, it is straightforward to consider more complicated coronal geometries with non-isotropic emission (Riaz et al., 2022). A ray-tracing code calculates how the corona (direct radiation from the corona) and the disk (thermal radiation of the disk returning to the disk) illuminate the disk (Mirzaev et al., 2024a). The latter is characterized by an electron density profile $n_{\rm e}=n_{\rm e}(r)$ (which can be specified by the user, but in the next section we will consider the simplest case of constant electron density) and its iron abundance $A_{\rm Fe}$. The ionization parameter of the disk $\xi$ at the radial coordinate $r$ is determined by the total incident photon flux $\Phi_{X}$ and the electron density at that radial coordinate
$\xi=\frac{4\pi\Phi_{X}}{n_{\rm e}}\,.$ (9)
The details of these calculations have been already presented in Mirzaev et al. (2024a) and have been already discussed in the literature; see, for example, Ingram et al. (2019), Dauser et al. (2022), and for a more pedagogical and detailed discussion Bambi (2024).
For the calculation of the reflection spectrum in the rest-frame of the material in the disk, we use the reflionx code (Ross & Fabian, 2005). We do not use the reflionx table because we want to calculate the reflection radiation with the correct spectrum of the incident radiation, while in the reflionx table it is assumed that the radiation illuminating the disk has a power law spectrum (which is a good approximation only for the direct radiation from the corona). After calculating the reflection spectrum from the direct radiation from the corona and the returning radiation of the thermal spectrum, the model includes the effect of the returning radiation of the reflection spectrum itself with a few iterations222ziji determines the final reflection spectrum in 4 iterations. As shown in Mirzaev et al. (2024a), there are only negligible differences if we go beyond the fourth iteration.. The details of these calculations have been already presented in Mirzaev et al. (2024a). At the end, we have a thermal component from the Novikov-Thorne accretion disk, a Comptonized component from the corona, and a reflection component calculated with the actual spectrum of the radiation illuminating the disk (direct radiation from the corona and returning radiation from the disk). Table 1 lists the parameters of the model in its simplest version.
@@ 3. Results
In this section, we present the predictions of ziji. We set the black hole mass to $10~{}M_{\odot}$. To maximize the impact of the returning radiation, we assume that the black hole spin parameter is $a_{*}=0.998$ and that the inner edge of the accretion disk is at the ISCO radius, $r_{\rm in}=r_{\rm ISCO}$. There is no emission from the plunging region $r<r_{\rm in}$. As shown in Fig. 1 in Mirzaev et al. (2024a), a significant fraction of the radiation emitted by the disk returns to the disk only in the very inner part of the accretion disk, when $r<2~{}r_{\rm g}$, where $r_{\rm g}=M$ is the black hole gravitational radius. Already at $r=2~{}r_{\rm g}$, most of the emitted radiation can escape to infinity and this fraction approaches 1 as the emission radius increases. We assume a lamppost corona and we set $h=5~{}r_{\rm g}$, where $h$ is the coronal height (i.e., the radial coordinate in Boyer-Lindquist coordinates of the point-like corona along the black hole spin axis).
@@
Figure 2.— Synthetic reflection spectra in the rest-frame of the material in the disk at different radial coordinates ($r=1.3$, 3, and 10 $r_{\rm g}$) when the Eddington-scaled luminosity of the source is $\lambda=0.01$. Yellow solid curves describe the direct incident spectra from the corona, black solid curves describe the incident spectra of the returning radiation of the thermal component, blue dashed curves are the reflection spectra produced by the direct radiation from the corona and the returning radiation of the thermal component, and red dashed curves are for the final reflection spectra including the returning radiation of the reflection radiation. In the top panels, the spectrum of the source is dominated by the spectrum from the corona ($\rho=L_{\rm corona}/L_{\rm disk}=100$). In the central panels, the luminosity of the corona and of the disk are equivalent ($\rho=1$). In the bottom panels, there is no corona and the reflection spectrum is entirely generated by the returning radiation of the thermal component ($\rho=0$). See the text for more details.
@@
The spectrum of the corona is a power law with an exponential high-energy cutoff $E_{\rm cut}$
$\frac{dN_{\rm c}}{dt_{\rm c}dE_{\rm c}}=K\,E_{\rm c}^{-\Gamma}\exp\left(-\frac{E_{\rm c}}{E_{\rm cut}}\right)\,,$ (10)
where the subindex ${\rm c}$ is used to indicate that a quantity is measured in the rest-frame of the corona, $K$ is a normalization constant, and we set the photon index $\Gamma$ to 2 and the high-energy cutoff $E_{\rm cut}$ to 300 keV. The total luminosity of the corona is
[…]
where we set $E_{\rm min}=1$ eV and $E_{\rm max}$ to infinity. Eq. (11) is used to evaluate $\rho=L_{\rm corona}/L_{\rm disk}$, which can be regulated by the normalization constant $K$. The spectrum of the corona observed far from the source is redshifted. The redshift factor is
$g=\left[\sqrt{-g_{tt}}\right]_{r=h,\theta=0}\approx 0.77\,,$ (12)
where $g_{tt}$ is evaluated at the point of the corona, where $h=5~{}r_{\rm g}$ in these simulations.
In Fig. 2, we show the spectra in the rest-frame of the material in the disk at the radial coordinates $r=1.3~{}r_{\rm g}$ (left panels), $3~{}r_{\rm g}$ (central panels), and $10~{}r_{\rm g}$ (right panels) when the Eddington-scaled luminosity of the source is $\lambda=0.01$. The top panels are for $\rho=100$ (the spectrum of the source is essentially the Comptonized spectrum from the corona), the central panels are for $\rho=1$ (the corona and the disk have the same luminosity), and the bottom panels are for $\rho=0$ (there is no corona and the reflection spectrum is generated exclusively by the returning radiation of the thermal spectrum of the disk). The electron density of the disk is $n_{\rm e}=10^{22}$ cm-3 at every radial coordinate and the ionization parameter $\xi$ is calculated from Eq. (9). The iron abundance is $A_{\rm Fe}=1$. In every panel, we show the incident spectrum of the direct radiation from the corona (yellow solid line), the incident spectrum of the returning radiation of the thermal component (black solid line), the reflection spectrum generated by these two components (blue dashed line), and the final reflection spectrum that includes even the returning radiation of the reflection spectrum (red dashed line).
Fig. 3 is like Fig. 2 but for $\lambda=0.1$. The higher luminosity of the source implies a higher ionization parameter and the spectra in the rest-frame of the material in the disk show fewer emission lines if compared to the panels in Fig. 2 with the same radial coordinate $r$. This is in particular the case for the panels for $r=1.3$ $r_{\rm g}$, because the ionization parameter is too high to have emission lines.
Fig. 4 shows the reflection spectra of the whole disk as observed far from the source. The radial profile of the ionization parameter $\xi$ in these simulations is shown in Fig. 5. If we considered a lower disk electron density $n_{\rm e}$, like the default value $10^{15}$ cm-3 in xillver, which is the most widely used value in current reflection studies, the inner part of the accretion disk would be fully ionized. We also note that $\rho$ is not the only parameter determining the relative contribution between the corona and the returning radiation of the thermal component in the generation of the reflection spectrum. The height of the corona $h$ also plays an important role. The impact of the coronal height on the reflection spectrum observed far from the source is shown in Fig. 6, where we assume $\rho=1$ and we only change the value of $h$.
Figure 4.— Synthetic reflection spectra of the whole disk as observed far from the source. The inclination angle of the disk with respect to the line of sight of the observer is $i=45^{\circ}$. Yellow solid curves describe the Comptonized spectra from the corona. Black solid curves describe the thermal spectra of the disk. Blue dashed curves describe the reflection spectra produced by the direct radiation from the corona and the returning radiation of the thermal spectrum of the disk (i.e., we do not include higher order disk reflection). Red dashed curves are for our final results for the reflection spectra (i.e., we include higher order disk reflection). In the top panels, the Eddington-scaled luminosity of the source is $\lambda=0.01$. In the bottom panels, we have $\lambda=0.1$. In the left panels, the luminosity of the source is dominated by the corona ($\rho=L_{\rm corona}/L_{\rm disk}=100$). In the right panels, there is no corona and therefore the reflection spectrum is produced by the returning radiation of the thermal component ($\rho=0$). In the central panels, the luminosity of the corona and of the disk are equivalent ($\rho=1$). See the text for more details.
Figure 5.— Radial profile of the ionization parameter $\xi$ in the simulations shown in Fig. 4. The Eddington-scaled luminosity of the source is $\lambda=0.01$ (left panel) and $\lambda=0.1$ (right panel). Dashed-green curves are for $\rho=L_{\rm corona}/L_{\rm disk}=0$, dashed-dotted-red curves are for $\rho=1$, and solid-blue curves are for $\rho=100$. The curves in the same style in the same panel refer to the ionization parameter calculated, respectively, without including (lower curves) and including (upper curves) the returning radiation of the reflection spectrum.
Figure 6.— As in Fig. 4 for $\rho=1$ and different values of the coronal height $h$. From left to right: $h=3$ $r_{\rm g}$, 5 $r_{\rm g}$, 10 $r_{\rm g}$, and 20 $r_{\rm g}$.